# ScaleRAFT: Cross-Scale Recurrent All-Pairs Field Transforms for 3D Motion Estimation

### v1

@@ VI Conclusion
In this article, we propose a new cross-scale recurrent all-pairs field transform for 3D motion estimation (ScaleRAFT), which extends the optical flow matching from the original image-to-image to image-to-scale-space, reducing the interference of scale changes on optical flow matching. Moreover, matching in scale space can also obtain depth motion clues, helping 3D flow methods break away from depth dependence and regression dependence. The comprehensive evaluation of different datasets shows that ScaleRAFT performs excellently in various 3D tasks, especially in areas with large scale changes. We hope this exciting result draws people to pay more attention to 3D motion estimation based on image feature matching.
VII Acknowledgement
We thank Professor Jiwen Lu of Tsinghua University for his constructive comments on this paper.
This research was supported by the National Natural Science Foundation of China (Grant nos. 61673220, 62106209,62372235), and the Project of SongShan Laboratory (Grant no. YYJC012022015).

### v2

@@ V Conclusion
In this article, we propose a new cross-scale recurrent all-pairs field transform for 3D motion estimation (ScaleFlow++), which extends the optical flow matching from the original image-to-image to image-to-scale-space, reducing the interference of scale changes on optical flow matching. Moreover, matching in scale space can also obtain depth motion clues, helping 3D flow methods break away from depth dependence and regression dependence.  We also propose a global iterative refinement (GIR) module and an initialization module to address the inherent problems of small network perception range and task alienation in iterative methods. These modules enable the network to perceive global motion and achieve more robust optimizer learning in fewer iterations. The comprehensive evaluation of different datasets shows that ScaleFlow++ performs excellently in various 3D tasks, especially in scenes with large scale changes (driving scenes). We hope this exciting result draws people to pay more attention to 3D motion estimation based on a monocular camera.